OpenStudy (anonymous):

sove the system by substitution 4x+6y=16 5x-6y=-88

OpenStudy (katrinakaif):

9x = -72 x = -8 Now subsitute that value back as 'x' and find y =)

OpenStudy (anonymous):

so the solution is (-72,-8)

OpenStudy (katrinakaif):

Hmm I don't see how y could be -8 though..

OpenStudy (anonymous):

(1) 4x+6y=16 (2) 5x-6y=-88 (1) x=4-(6/4)y divide by 4 to isolate x (2) 5[4-(6/4)y] -6y=-88 substitute x from 1st equation into 2nd equation simplify: 20-(30/4)y -(24/4)y=-88 -(54/4)y=-88-20 -13.5y=-108 y=8 substitute y=8 into either original equation to solve for x Equation 1) 4x+6(8)=16 simplify and solve for x 4x+48=16 4x=16-48 x=-32/4 x=-8 solution set:(-8, 8)

OpenStudy (precal):

4x+6y=16 4(-8)+6y=16 -32+6y=16 add 32 to both sides 6y=48 divide 6 to both sides y=8